OpenStudy (anonymous):

The area of a rectangle is the product of width and length. If a rectangle has length 15m sqrt 13m and width sqrt 9m, which of the following is the area of the rectangle? 135m2 sqrt 13 135m sqrt 13m 45m2 sqrt 13 45m sqrt 13m

OpenStudy (anonymous):

help please

OpenStudy (anonymous):

so it has length \[15m \sqrt{13m}\] and width \[\sqrt{9}\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

yes @JamesWolf

OpenStudy (anonymous):

what are you struggling with, the formula or the operation of multiplying it all?

OpenStudy (anonymous):

multiplication of it all, what to do with the radical?

OpenStudy (anonymous):

leave it, your answers are in the form of (some number) x (some radical)

OpenStudy (anonymous):

so for example \[25\sqrt{42} \times 4 = 100\sqrt{42}\]

OpenStudy (anonymous):

so 15m sqrt 117m?

OpenStudy (anonymous):

15m \[\sqrt{117m}\]

OpenStudy (anonymous):

well you can actually square root the 9. \[\sqrt{9} = 3\]

OpenStudy (anonymous):

which gives you \[15m \sqrt{13m} \times 3m = area\] =

OpenStudy (anonymous):

so 45m \[\sqrt{13m}\] ?

OpenStudy (anonymous):

or is something squared?

OpenStudy (anonymous):

yeah right on! but remember that your timesing the "m" too so you get \[45 m^{2}\sqrt{13m}\]

OpenStudy (anonymous):

and im almost certain "timesing" isnt a word

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

no problem :)